### v1

# Right for Right Reasons: Large Language Models for Verifiable Commonsense Knowledge Graph Question Answering

Abstract
Knowledge Graph Question Answering (KGQA) methods seek to answer Natural Language questions using the relational information stored in Knowledge Graphs (KGs). With the recent advancements of Large Language Models (LLMs) and their remarkable reasoning abilities, there is a growing trend to leverage them for KGQA. However, existing methodologies have only focused on answering factual questions, e.g., "In which city was Silvio Berlusconi’s first wife born?", leaving questions involving commonsense reasoning that real-world users may pose more often, e.g., "Do I need separate visas to see the Venus of Willendorf and attend the Olympics this summer?" unaddressed. In this work, we first observe that existing LLM-based methods for KGQA struggle with hallucination on such questions, especially on queries targeting long-tail entities (e.g., non-mainstream and recent entities), thus hindering their applicability in real-world applications especially since their reasoning processes are not easily verifiable. In response, we propose Right for Right Reasons ($R^{3}$), a commonsense KGQA methodology that allows for a verifiable reasoning procedure by axiomatically surfacing intrinsic commonsense knowledge of LLMs and grounding every factual reasoning step on KG triples. Through experimental evaluations across three different tasks—question answering, claim verification, and preference matching—our findings showcase $R^{3}$ as a superior approach, outperforming existing methodologies and notably reducing instances of hallucination and reasoning errors.
1 Introduction
Knowledge Graphs (KGs) have been widely used as a structured format for storing and representing relational information. Efficiently querying KGs to obtain the required knowledge is a long-standing problem, for which query languages such as RQL Karvounarakis et al. (2002) and SPARQL Prud’hommeaux and Seaborne (2008) have been developed. However, writing queries in these languages requires expertise which limits the accessibility of KGs to inexpert users. Knowledge Graph Question Answering (KGQA) Zheng et al. (2017); Berant et al. (2013a); Yih et al. (2016) is an established research field that facilitates access to KGs by providing factual answers to natural language (NL) questions using KGs.
Recently, the promising performance of Large Language Models (LLMs) in reasoning-related tasks has encouraged their application in KGQA research Baek et al. (2023); Guan et al. (2023b); Li et al. (2023a). While these works have significantly enhanced the performance of KGQA systems, their primary focus has been on addressing factoid questions, such as "In which city was Silvio Berlusconi’s first wife born?", which can be answered using only the knowledge graph (KG) facts. However, real-world user queries often extend beyond the factoid knowledge stored in the KG. For example, answering a question such as "Do I need separate visas to see the Venus of Willendorf and attend the Olympics this summer?" requires both KG triples indicating the locations of Venus of Willendorf and the place where this summer’s Olympics is taking place, as well as commonsense reasoning about how one can identify whether traveling to those countries requires separate visas or not.
Commonsense reasoning is one of the most significant capabilities offered by LLMs Shen and Kejriwal (2023a). Therefore, it may seem straightforward to leverage the LLMs to reason over a set of retrieved KG facts to perform commonsense KGQA. However, LLMs are still susceptible to introducing ungrounded or incorrect information to their reasoning process – a phenomenon called hallucination Ye et al. (2023); Tonmoy et al. (2024). In conducting commonsense KGQA, LLMs may exhibit hallucinations both by introducing ungrounded factual information as well as making incorrect commonsense inferences. Hence, verifiability of the reasoning process is crucial to ensure the reliability of the final answer, especially in high-stake applications. Regrettably, none of the existing LLM-enhanced KGQA methodologies answer queries following a verifiable scheme.
In this paper, we experimentally show that the performance of existing KGQA methods is critically hindered by the hallucination issue when faced with questions involving commonsense reasoning. This issue is particularly exacerbated for questions about long-tail knowledge, i.e., questions targeting obscure or recent entities, and personalized questions. To address this challenge, we introduce Right for Right Reasons $(R^{3})$, a verifiable methodology for performing KGQA using LLMs. $R^{3}$ makes both aspects of commonsense KGQA reasoning, factoid steps and commonsense inferences, verifiable. For the commonsense inference aspect, it axiomatically surfaces the commonsense knowledge required for answering the question that is intrinsic to the LLM parameters. Also, it casts the KGQA task into a tree-structured search in which all factual reasoning steps are enforced to be grounded on a subset of the relevant KG triples which enables the verification of factual reasoning steps. We compare $R^{3}$ against current LLM-based KGQA methodologies and pure LLM methods on three different tasks: question answering, claim verification, and KG-based preference matching. The results demonstrate that $R^{3}$ leads to a considerable reduction in hallucination and reasoning errors while often improving accuracy and offering robustness to entity popularity.
2 Background
2.1 Reasoning with Large Language Models
Despite being originally designed for text generation, LLMs have shown outstanding performance when applied to several other NLP sub-fields Chang et al. (2023). Particularly, the reasoning capability of LLMs has attracted considerable interest in AI research Arora et al. (2022); Sun et al. (2022). Several works have studied different reasoning skills of LLMs such as arithmetic reasoning Yuan et al. (2023), logical reasoning Liu et al. (2023), and commonsense reasoning Bian et al. (2023); Shen and Kejriwal (2023b). These abilities make LLMs apt candidates for being used as a reasoner in tasks from other domains Ren et al. (2023); Song et al. (2023); Clusmann et al. (2023).
2.2 Commonsense Question Answering
The general knowledge and conception about the world that humans possess, and their ability to reason about it is called commonsense reasoning and is a crucial cognitive ability of humans. It is also an important reasoning skill based on which AI agents are evaluated Liu et al. (2021); bauer-etal-2022-nlp; Wang et al. (2023). LLMs have shown outstanding commonsense reasoning skills and have narrowed the gap between them and humans on available datasets Guan et al. (2023a); Bian et al. (2023). Most existing datasets such as CommonsenseQA Talmor et al. (2018) and PhysicalQA Bisk et al. (2020) contain questions about concepts rather than entities. Recently, StrategyQA Geva et al. (2021a) and Creak Onoe et al. (2021b) have been proposed as datasets for commonsense reasoning about entities that can be used to introduce commonsense reasoning to KGQA.
2.3 Knowledge Graph Query Answering
The overall objective of KGQA is to answer an NL query using the KG facts. Most existing works on KGQA focus on converting the NL query into a structured formal query in a language such as SPARQL, executing the query to retrieve the required knowledge, and finally reasoning over the retrieved facts to obtain the final answer. This idea, referred to as semantic parsing Lan et al. (2021); Gu and Su (2022); Cheng et al. (2022), often involves the data and computationally expensive process of fine-tuning with thousands of labeled examples Chen et al. (2021); Shu et al. (2022). Recently, KB-BINDER has suggested a training-free semantic parsing methodology using the in-context learning ability of LLMs with few-shot examples Li et al. (2023b). Novel LLM-based methods beyond semantic parsing approach have also been proposed. KAPING Baek et al. (2023) introduced an efficient LLM-enhanced KGQA model that finds the relevant sub-graph to the query via dense retrieval and uses the LLM to reason over it in a zero-shot manner. KGR Guan et al. (2023b) proposed the idea of allowing LLMs to make claims, retrofitting those claims on the KG facts, and finally reasoning using the corrected claims. However, all existing works on KGQA are designed to answer factoid queries, and none of them has considered queries involving commonsense reasoning.
3 Methodology
3.1 Problem Formulation
In this paper, we propose a methodology for performing commonsense KGQA that is easily extended to other related tasks such as claim verification and KG-based preference matching. The input to the problem is a NL sentence posed by the user that can be either a question in the form of an interrogative sentence, or a claim or need expressed as an imperative sentence. We use the term query, denoted by $q$, to refer to the input in all cases. The query mentions a set of anchor entities $\mathcal{E}^{q}$. A KG $K=(\mathcal{E},\mathcal{R})$ is assumed to be given, where $\mathcal{E}$ and $\mathcal{R}$ denote its set of entities and relations respectively, such that $\mathcal{E}^{q}\subset\mathcal{E}$. The objective is to follow a sequence of reasoning steps $\mathcal{S}^{q}$ to find $a^{q}\in\mathcal{O}^{q}$, the answer to the query, such that verifying the correctness of every $s^{q}_{i}\in\mathcal{S}^{q}$ is possible. Here, $\mathcal{O}^{q}$ denotes the set of possible options.
Figure 1: An example workflow of commonsense KGQA procedure using $R^{3}$.
3.2 Right for Right Reasons
Our proposed methodology casts the problem of commonsense KGQA as a tree-structured search, in which every reasoning step is either grounded on KG facts, or based on surfaced commonsense axioms, a key property that makes the reasoning procedure completely verifiable. The overall workflow of $R^{3}$ for answering a query is shown in Figure 1. In brief, $R^{3}$ first identifies the anchor entities of a query and obtains the relevant sub-graph for these entities. Next, it surfaces a commonsense axiom from the LLM that will guide the reasoning steps in that branch of the search tree. Then, at each depth level of the tree, it checks whether the commonsense axiom can be satisfied with the available KG facts, and if possible, provides an answer grounded on a subset of them. If the available KG triples are insufficient, by backward-chaining from the axiom, it selects the next entity to obtain its relevant KG sub-graph to continue the search. Each branch can continue up to a maximum depth, and if an answer is not obtained at its bottom, a new commonsense axiom will be surfaced which will guide search in a new branch until the search tree reaches its maximum breadth. Here, we explain each component of $R^{3}$ in more detail.
3.2.1 Obtaining Relevant Sub-graph
The query answering process begins by extracting $\mathcal{E}^{q}$ from $q$. Most existing works perform this extraction using entity linking techniques Li et al. (2020); Ayoola et al. (2022). However, since existing entity linkers may fail to extract recent or obscure entities from the query, we also use an LLM-based module with few-shot examples to obtain another set of entity names, and consider the union of the two sets as the final set of entities. Formally,
$\mathcal{E}^{q}=\mathtt{EL}(q,\mathcal{K})\cup\mathtt{LLM_{E}}(q),$ (1)
where $\mathtt{EL}$ is an entity linker module and $\mathtt{LLM_{E}}$ is the LLM-based module that identifies anchor entities mentioned in $q$. Once the anchor entities are identified, we extract $\mathcal{K}^{q}\subset\mathcal{K}$, the sub-graph of $\mathcal{K}$ within the 1-hop neighborhood of $\mathcal{E}_{q}$.
$\mathcal{K}^{q}=\{(h,r,t)|(h,r,t)\in\mathcal{K}\land h\in\mathcal{E}^{q}\}.$ (2)
3.2.2 Surfacing Commonsense Axioms
The commonsense knowledge that LLMs have obtained during their training process is intrinsic to their parameters, and they can use it to answer queries given a set of retrieved facts. Existing LLM-based methods that are designed for tackling the factoid KGQA problem can approach commonsense KGQA using this intrinsic capability of their LLM component. However, since the set of commonsense axioms the reasoner has used is not known, the reasoning process is not verifiable. To address this issue, $R^{3}$ axiomatically surfaces this intrinsic knowledge of the reasoner and uses it to guide the reasoning process. In other words, its reasoner is enforced to state the premises required for concluding an answer as a set of atomic factoid clauses and try to find the answer by identifying whether those clauses are satisfied when their variables are grounded on the KG entities, and their predicates and functions on KG relations. For example, when given a query "Would it make sense for Virginia Raggi to ask for a quinceañera?", the reasoner surfaces the axiom: "If Virginia Raggi is a girl from Latin America and her age is near 15, it would make sense for her to ask for a quinceañera."
Formally, given $\mathcal{E}^{q}=\{e_{1}^{q},...,e^{q}_{|\mathcal{E}_{q}|}\}$, a commonsense axiom $I_{q}$ is an NL representation of the First-Order Logic (FOL) expression
[…]
3.2.3 Sub-graph Pruning
Once a commonsense axiom is surfaced, $R^{3}$ tries to identify the satisfiability of the premises based on the KG triples. Since the number of triples in $\mathcal{K}^{q}$ may be large, we need to first prune the set of available KG triples. To this end, as in  Baek et al. (2023), we use off-the-shelf dense retrievers Song et al. (2020); Karpukhin et al. (2020); Xiong et al. (2020) to obtain $\mathcal{T}^{q}_{i}\subset\mathcal{K}^{q}$, the subset of triples that have the most semantic similarity to the commonsense axiom $I^{q}_{i}$. Since filtering triples by only considering semantic similarity may lead to a high risk of losing some useful triples, we also use an LLM module with few-shot examples to pick relevant triples to the axiom from a subset of the sub-graph triples to reduce the chance of this information loss. Formally we have the Sub-graph Pruning module $\mathtt{SGP}$ as
$\begin{split}&\mathtt{SGP}({I}^{q}_{i},\mathcal{K}^{q})=\\ &\mathtt{top-k}_{t\in\mathcal{K}_{q}}(sim(\boldsymbol{t},\boldsymbol{I}^{q}_{i}))\cup\mathtt{LLM}_{T}(\mathcal{K}^{q},{I}^{q}_{i}),\\ &\mathcal{T}^{q}_{i}=\mathtt{SGP}({I}^{q}_{i},\mathcal{K}^{q}),\end{split}$ (4)
in which $sim$ denotes the Euclidean similarity between $\boldsymbol{t}$ and $\boldsymbol{I}^{q}_{i}$, the embedding vectors of the triple $t$ and the axiom $I^{q}_{i}$, $\mathtt{top-}k$ operator returns the first $k$ elements of the sorted list of triples by their similarity score in descending order, and $\mathtt{LLM}_{T}$ is an LLM-based module that returns a subset of $\mathcal{K}^{q}$ that are relevant to $I^{q}_{i}$.
3.2.4 Fact-Grounded Answer Selection
[…]
$a^{q}=\mathtt{answer}(q,\boldsymbol{I}^{q}_{i},\mathcal{T}^{q}_{i}),$ (5)
where $\mathtt{answer}$ is the LLM-based module determining the final answer.
3.2.5 Missing Evidence Identification
The set of retrieved facts may be insufficient in two cases: either the query targets a different entity, as in multi-hop questions, or the facts required for grounding at least one premise were mistakenly pruned. In this step, the reasoner is asked to consider the set of unsatisfied premises and the existing facts to first identify what additional evidence must be obtained that is currently missing. Then, it has to identify the anchor entity $e_{m}$ that its triples can provide the missing information. If the anchor entity is already in, $\mathcal{E}_{q}$, the next top $k$ relevant facts about it will be picked for the next step. Otherwise, the reasoner is asked to propose the next entity and extract its name from $\mathcal{K}^{q}$. The next entity is then added to $\mathcal{E}_{q}$, and the process of sub-graph extraction and pruning is executed for it. Formally,
$\begin{split}e^{m}&=\texttt{MEI}(q,I^{q}_{i},\mathcal{T}_{i}),\\ \mathcal{E}^{q}_{j+1}&=\mathcal{E}^{q}_{j}\cup\{e^{m}\},\\ \mathcal{K}^{q}_{j+1}&=\mathcal{K}^{q}_{j}\cup\{(h,r,t)|(h,r,t)\in\mathcal{K}\land h\in e^{m}\},\\ \mathcal{T}_{i}&=\mathcal{T}_{i}\cup\mathtt{SGP}(\mathcal{K}^{q}_{j+1},I^{q}_{i}),\\ \end{split}$ (6)
where MEI is the module identifying entity $e^{m}$. This procedure continues until an answer is found or the maximum depth is reached for the branch. In case the maximum depth for a branch is reached without obtaining an answer, a new commonsense axiom will be generated to form a new branch.
3.3 Comparison to Existing KGQA Methods
$R^{3}$ is the first KGQA approach that supports commonsense queries in a verifiable manner, since every factual reasoning step is grounded on particular KG triples, and its commonsense reasoning assumptions are surfaced in the form of axioms. Although KGR Guan et al. (2023b) retrofits its factual claims on the KG, its commonsense reasoning process is implicit. Semantic parsing methods are only designed for factoid queries and cannot address commonsense queries. Finally, KAPING Baek et al. (2023), despite its strong performance on single-hop factoid queries, cannot answer multi-hop questions because it has no particular mechanism for traversing the KG. A summary of key properties of existing KGQA methods and their comparison to $R^{3}$ is provided in Table 1.
Method Factoid QA Verifiability Commonsense No training Multi-hop
Classical Semantic Parsing
KB-BINDER
KAPING
KGR
$R^{3}$
Table 1: Comparison of $R^{3}$ properties against existing KGQA Methods
4 Experiments
We empirically evaluate $R^{3}$ on three tasks: Question answering, claim verification, and KG-based preference matching. All tasks are closely related to KGQA and involve commonsense reasoning.
4.1 Task Description
Question Answering. In this task, a question that is formed around one or multiple KG entities and requires commonsense reasoning is provided. The reasoner is required to find the answer, which is either "Yes" or No.
Claim Verification Claim verification is very similar to question answering. Here, an imperative sentence making a claim about a number of entities is provided. The reasoner has to use the KG information to decide whether the claim is "Correct" or "Incorrect".
KG-based Preference Matching In this task, a NL query has been provided that explains the user’s requirement. Also, a personal KG containing evidence about the user’s preferences and restrictions is presented to the reasoner. The reasoner has to choose the item that matches both the user’s query and their personal restrictions.
4.2 Datasets
Due to the lack of existing datasets, we modify three existing datasets to make them suitable for our tasks and make them publicly available to encourage research on commonsense KGQA. Examples of these modifications are shown in Table 3.
Task Model Accuracy FActScore Reasoning
Original Long-Tail Original Long-Tail Original Long-Tail
Question Answering 0-shot CoT 0.70 0.32 0.63 0.54 0.90 0.89
2-shot CoT 0.70 0.43 0.64 0.52 0.92 0.90
KAPING 0.72 0.67 0.74 0.59 0.86 0.83
KB-BINDER 0.11 0.08 - - - -
KGR 0.39 0.13 0.61 0.47 0.70 0.65
$R^{3}$ 0.82 0.69 0.97 0.96 0.97 0.95
Claim Verification 0-shot CoT 0.89 0.35 0.76 0.59 0.93 0.91
2-shot CoT 0.92 0.41 0.78 0.58 0.93 0.92
KAPING 0.91 0.81 0.81 0.75 0.90 0.88
KB-BINDER 0.35 0.14 - - - -
KGR 0.80 0.20 0.70 0.58 0.74 0.71
$R^{3}$ 0.85 0.85 0.98 0.98 0.97 0.96
Table 2: Results for all methods on the question answering and claim verification task on both the original and modified (long-tail) queries. FaCTScore and Reasoning are human evaluated metrics.
Question Answering Early KGQA datasets consisted of simple questions that can be answered using a single KG triple. Recently, datasets containing more complex questions by introducing multi-hop reasoning have been proposed Berant et al. (2013b); Trivedi et al. (2017); Gu et al. (2021). However, all KGQA datasets contain factoid questions, which do not require commonsense reasoning to answer. Some datasets exclusively focus on evaluating commonsense reasoning Talmor et al. (2018); Boratko et al. (2020); Sap et al. (2019), but their questions target concepts (e.g., river, mountain, etc.) rather than KG entities (e.g., specific people, locations, etc.).
To overcome this challenge we modify StrategyQA Geva et al. (2021b), a QA dataset with Yes/No questions that target entities from Wikipedia [^1] https://www.wikipedia.org/ articles. We select a subset of 150 questions for which the required factual knowledge for answering them is present in Wikidata [^2] https://www.wikidata.org/ or that can be rewritten as such queries by targeting them on new entities. The questions mostly target famous entities that LLMs can answer using their internal knowledge without hallucinating or even needing a KG. Since we are particularly interested in studying the hallucination behavior of LLM-based KGQA methods on long-tail knowledge, for each query, we also write a counterpart targeting long-tail knowledge by substituting its entities with obscure entities of the same types. We use the number of Wikidata triples and Google Search results as measures of popularity.
Claim Verification For the task of KG-based claim verification, we use Creak Onoe et al. (2021a), a dataset containing True/False claims written by crowd workers using Wikipedia. We follow a similar procedure applied to the Question Answering dataset to select 150 claims and write their long-tail counterparts.
KG-based Preference Matching Recipe-MPR Zhang et al. (2023) is a preference matching dataset that contains NL queries expressing a user’s preference toward recipes and often targeting multiple aspects. The reasoner has to choose the recipe that satisfies all aspects among five options. The multi-aspect nature of its queries and the necessity for performing logical reasoning make it a relevant dataset to our work. However, its queries are not personalized, meaning that the correct recipe does not require reasoning over the user’s preferences and restrictions beyond those stated in the query. In real-world applications, the "correct" item is different for each user considering their personal preferences and restrictions. To bridge this gap, we first extract 100 queries from Recipe-MPR dataset that require commonsense reasoning and add a synthetic personal KG for the user posing the query. We also add a sixth option that matches every preference aspect of the query but violates at least one personal requirement that can be inferred from the user’s personal KG.
Question Answering
Exemplar Original Question
Did Alan Turing suffer from the same fate as Abraham Lincoln?
Exemplar Modified Question
Did Ivan Shuisky suffer from the same fate as Benny Frey?
Claim Verification
Exemplar Original Claim
The Bugs Bunny cartoons were influenced by the cartoon Rick and Morty.
Exemplar Modified Claim
Giovanni Battista Casti’s works may be influenced by Maria Grazia Lenisa’s poems.
Preference Matching
Exemplar Query
Sam: I like eating pulled meats, but not beef or chicken.
Original Options
[…]
Added Option
[…]
Personal KG
(Sam, occupation, painter), (Sam, age, 29), (Sam, medical condition, allium allergy), …
, (Sam, religion, Christianity), (Sam, medical condition, lactose intolerance).
Table 3: Exemplar queries form Datasets used for each task and modifications applied to them. Modified queries in Question answering and Claim verification target obscure entities to evaluate robustness to popularity shift. The synthetic KG and the new option add personalization aspect to the Preference Matching task.
4.3 Experimental Setup
We compare $R^{3}$ against LLM baselines with Chain-of-Thought (CoT) prompting, both in zero-shot Kojima et al. (2022) and few-shot ($k=2$) settings  Wei et al. (2022) to evaluate the need for a KG to answering these queries, and three recent LLM-enhanced KGQA models, KB-BINDER Li et al. (2023a), KGR Guan et al. (2023b), and KAPING Baek et al. (2023). For question answering and claim verification tasks, we evaluate all methods on both original queries (targeting famous entities) and modified queries (targeting long-tail entities) to study their robustness to popularity shift. We use GPT-3.5 Turbo as the LLM for all models. In addition to accuracy, we perform human evaluation to measure factual and reasoning faithfulness. In particular, we use FActScore Min et al. (2023), which measures the percentage of atomic facts in an LLM’s response that is supported by a knowledge base, and Reasoning score, which measures the proportion of LLM responses in which there are no logical reasoning errors. For KG-based preference matching, our human evaluation consists of measuring Accuracy of Reasons which is simply the fraction of correct answers that were obtained from correct reasons.
4.4 Results
4.4.1 Question Answering
The results for the question answering task are presented in table 2. $R^{3}$ beats all baselines, achieving an accuracy of 0.82 and 0.69 in the original and long-tail settings respectively. Although the strongest baseline, KAPING, achieves comparable accuracy, human evaluation reveals that KAPING’s answers are far less reliable than those of $R^{3}$. KB-BINDER’s performance is much lower than other methods, because KB-BINDER is a semantic parsing method that only supports factoid queries and not ones that require commonsense reasoning. Although 0-shot and few-shot CoT achieve 0.70 accuracy on the original queries, their accuracies drop significantly in the long-tail setting to 0.32 and 0.43 respectively. We also observe in the long-tail setting a sharp increase in the number of questions for which the LLM responds “I don’t know.”
Among all methods, $R^{3}$ hallucinates the least, with the highest FActScores, 0.97 and 0.96, in the original and long-tail settings respectively. KAPING’s FActScores, 0.74 and 0.59, are significantly lower than $R^{3}$, despite leveraging dense retrieval. This is because KAPING’s retrieval is limited to entities in the question, which works only for single-hop queries. For multihop queries, KAPING resorts to the LLM’s internal knowledge. From our LLM baselines, we observe low FActScores, indicating that LLM’s internal knowledge is insufficient. In contrast, $R^{3}$ enforces strict grounding on the KG for reasoning, and has an iterative mechanism for identifying what additional facts are required, which leads to near perfect FActScores.
Not only are the FActScores of baseline methods significantly lower than $R^{3}$, but we also observe for all baselines a significant decrease in FActScore on long-tail queries. For instance, KAPING’s FActScore drops by 0.15 from 0.74 to 0.59. These results show that baseline method hallucinations are exacerbated in the long-tail setting due to LLMs being unable to faithfully recall long-tail knowledge. For KAPING, we also observe that the entity linker fails more often to identify long-tail entities, which inevitably leads to ungrounded hallucinated answers in the absence of relevant triples. In contrast, $R^{3}$ maintains a high FActScore in both the original and long-tail settings with respective scores of 0.97 and 0.96, which indicate its robustness to shifts in entity popularity.
$R^{3}$ also maintains the highest reasoning score compared to all baselines, achieving a score of 0.97 and 0.95 in the original and long-tail settings respectively, compared to the next best method, few-shot CoT, which achieves reasoning scores of 0.92 and 0.90. Because $R^{3}$ makes the commonsense inference process explicit by axiomatically surfacing the commonsense inference rules, $R^{3}$ provides both more verifiable and faithful chains of reasoning with less errors. In contrast, KAPING has a low reasoning score. We qualitatively observe that due to the low precision of the facts retrieved by KAPING, the LLM is frequently misled by the irrelevant facts. Elsewhere, KGR has the lowest reasoning score. Without CoT, KGR’s initial response often contains poor reasoning, which then leads to poor retrofitting and thus a low FActScore as well. Note that we do not perform human evaluation for KB-BINDER since it is a semantic parsing method that outputs SPARQL queries which are incompatible with FActScore and reasoning scores.
4.4.2 Claim Verification
The results for the claim verification task are presented in table 2. Although 2-shot CoT beats our method on the original queries, our method is robust in the long-tail setting, achieving the same accuracy as the original setting whereas 2-shot COT’s accuracy drops significantly by 0.51.
We observe that again $R^{3}$ maintains the highest FActScore, 0.98, in both the original and long-tail settings. In contrast, similar to the question answering task, all baseline methods have significantly lower FActScores that also decrease significantly in the long-tail setting. The low and decreasing FActScores in both the question answering and claim verification task crucially demonstrate that LLMs suffer from high rates of hallucination which are exacerbated in long-tail settings.
$R^{3}$ also maintains the highest reasoning score among all methods, 0.04 better than the next-best method which is few-shot CoT. Interestingly, with few-shot CoT, we qualitatively observe that the LLM at times erroneously follows the reasoning strategies in the examples. We believe that explicitly surfacing commonsense axioms is crucial for correctly guiding the subsequent reasoning. Again, KAPING’s low precision KG retrieval misleads the LLM, resulting in low reasoning scores, and KGR’s poor reasoning leads to suboptimal initial responses that KGR has difficulty retrofitting.
4.4.3 Preference Matching
Results of the preference matching task are provided in Table 4. Since the personal KG does not support SPARQL queries, KB-BINDER cannot be evaluated on it. KGR and pure LLM baselines also cannot be evaluated on this task since they can only make claims or provide answers about entities that LLMs are aware of, and not about users in a synthetic dataset. So, the only relevant baseline is KAPING. Results of this comparison vividly identify that on the challenging task of personalized preference matching, $R^{3}$ obtains a considerably higher accuracy. We also observe that the Accuracy of reasons for $R^{3}$ is more than double the number for KAPING, which again reflects its stronger commonsense reasoning ability due to its special approach for surfacing commonsense axioms.
Method Accuracy Accuracy of Reasons
KAPING 44 31.8
$R^{3}$ 57 70.17
Table 4: Results of Accuracy and Accuracy of reasons (%) for preference matching task
5 Conclusion
We proposed $R^{3}$, a novel framework that enables answering KG queries involving commonsense reasoning using LLMs in a verifiable manner by axiomatically surfacing their intrinsic commonsense knowledge. Key experimental results exhibit the efficacy of $R^{3}$ across different tasks related to KGQA and its superior performance to existing baselines. The promising performance of $R^{3}$ combined with its verifiability and robustness to entity popularity opens up possibilities for versatile future extension to address broader ranges of tasks and improve the flexibility and accessibility of KGs and reliability of LLM-based reasoners.
6 Limitations
In this section, we honestly and faithfully present the limitations of our current research and briefly discuss the possible future directions. The quality of the reasoning process in our methodology relies on the quality of the axioms generated. We observe through our experiments that in cases where the quality of axioms is insufficient, the reasoner is misled resulting in an undetermined answer at the end of the exploration budget identified. It is necessary to mention that in our experiments, we used $b=2$, a relatively small value for the maximum breadth since we expected the model to obtain an answer in a rather small number of efforts. Furthermore, due to the importance of avoiding hallucination, our model takes a conservative and rigorous approach to ground every factual premise on KG triples. Therefore, our model typically leaves more questions unanswered than other baselines. Obviously, we considered the unanswered questions as an incorrect response in calculating the accuracy.
Furthermore, as in most LLM-based models, for having a proper performance, LLM-based components of our model require clear explanation of the task provided in the prompts to them, as well as a number of few-shot examples that clarify the intent of the task description further.
We consider further studies into the above limitations as open areas of future work. Studying the trade-off between rigorousness and the rate of answering, as well as studying the robustness of our model to different prompting styles are key research questions that we consider for future.